OpenStudy (anonymous):

a 727 jet needs to attain a speed of 200 mph to take off. if it can accelerate from 0 to 200 mph in 30 seconds, how long must the runway be? (assume constant acceleration)

OpenStudy (ash2326):

v-u=at v=200mph =89.408 m/s u=0 t=30 a=? v-u=at 89.408=a*30 a=2.9802m/s^2 now the runway's length be s v^2-u^2=2as 89.408^2=2*2.9802*s s=89.408^2/(2*2.9802) s=1341.15 meters in miles 0.833